OpenStudy (anonymous):

Just another cute problem, Prove that: \( \large \tan \frac \pi {16} + 2 \tan \frac \pi {8} + 4 = \cot \frac{\pi}{16} \) Enjoy!

OpenStudy (anonymous):

If we substitute cos and sin instead of tan and then use the trig identity that\[\sin^{2}x + \cos^{2}x = 1\]we can prove it, I think. But I may have commited a minor mistake in my proof. I will proofread it.

OpenStudy (dumbcow):

hmm since there is no variable, we are only proving one specific case so verifying rhs equals lhs is sufficient no?

OpenStudy (anonymous):

The quickest way seems to be rewritting sin and cos using series expasions and to do long division on them to solve it, haha. :-)

OpenStudy (anonymous):

@dumbcow That works! as long as you are deriving RHS wholly from the LHS. PS: The actual problem came with 4 options.

OpenStudy (anonymous):

math is not cute. math is ugly. very ugly.

OpenStudy (anonymous):

I derived it from half angles formulas. I think it's an okay proof, but too messy. I would like to see a cleaner and more elegant proof than mine, if possible :-)

OpenStudy (anonymous):

My proof is of 3 lines.

OpenStudy (anonymous):

I figured it out that my proof is very inelegant. Still, any tips for the cuter solution? :-)

OpenStudy (anonymous):

Use the double angle formula for tan to get \[ \tan(\frac \pi 8)=\sqrt{2}-1 \] Do the same trick to find \[\tan(\frac \pi {16})=-1-\sqrt{2}+\sqrt{4+2 \sqrt{2}} \] From now it is easy. I am sure you have an easier proof.

OpenStudy (anonymous):

tan (pi/16)-cot(pi/16)+2tan(pi/8)+4 =(2tan(pi/16)/tan(pi/8)) / tan(pi/16) +2 tan(pi/8) +4 =-2cot(pi/8)+2tan(pi/8) +4 =-2(cos pi/8 / sin pi/8 + sinpi/8 /cos pi/8) +4 =-2sin pi/4 / (1/2 sin pi/4) +4 =0

OpenStudy (anonymous):

\[\cot (\frac \pi {16}) =1+\sqrt{2}+\sqrt{2 \left(2+\sqrt{2}\right)} \] so \[-1-\sqrt{2}+\sqrt{4+2 \sqrt{2}} + 2 + 2( \sqrt{2}-1) +4=1+\sqrt{2}+\sqrt{2 \left(2+\sqrt{2}\right)}=\cot(\frac \pi{16}) \]

OpenStudy (anonymous):

Use the formulas \[\tan (\frac x 2) =\frac{ 1- \cos(x)}{\sin(x)}\\ \cot (\frac x 2) =\frac{ 1+ \cos(x)}{\sin(x)}\\ \] Let \[ x= \frac \pi 8 \] With this notation we have to prove \[ \tan(\frac x 2) + 2 \tan(x) - \cot(\frac x 2)=-4\\ \tan(\frac x 2) + 2 \tan(x) - \cot(\frac x 2)=\\ \frac{ 1- \cos(x)}{\sin(x)} + 2 \frac {\sin(x)}{\cos(x)}- \frac{ 1+ \cos(x)}{\sin(x)}=\\ -2 \frac {\cos(x)}{\sin(x)}+ 2 \frac {\sin(x)}{\cos(x)}=\\ \frac {-2\left( \cos^2(x) -\sin^2(x)\right)} {\sin(x) \cos(x)}=\\ \frac { -2 \cos(2x)}{\frac 12 \sin(2 x)}\\ -4 \frac { \cos(2x)}{ \sin(2 x)}= -4\\ \] Since \[ 2 x =\frac \pi 4\]

OpenStudy (anonymous):

An easier proof. It easy to show that \[ \tan(v) - \cot (v) =- 2 \cot(2v) \] Take \[ v=\frac \pi{16} \] We get \[\tan(\frac \pi{16}) - \cot (\frac \pi{16}) =- 2 \cot(\frac \pi{8}) \] What we have to prove becomes \[ - 2 \cot(\frac \pi{8}) + 2 \tan(\frac \pi{8}) =-4\\ -\cot(\frac \pi{8}) + \tan(\frac \pi{8})=-2\\ \text { By the first formula }\\ - 2 \cot(\frac \pi 4) =-2\\ -2(1) =-2 \]

OpenStudy (anonymous):

@bmp: I used the identity \(\tan A = \cot A -2\cot 2A \)

OpenStudy (anonymous):

Indeed, that's simpler than what I've done. I did it very much like the first solution posted by @eliassaab.